A fitness course claims that it can improve an individual's physical ability. To test the effect of a physical fitness course on one's physical ability, the number of sit-ups that a person could do in one minute, both before and after the course, was recorded. Ten individuals are randomly selected to participate in the course. The results are displayed in the following table. Can it be concluded, from the data, that participation in the physical fitness course resulted in significant improvement?

Let d=(number of sit-ups that can be done after taking the course)−(number of sit-ups that can be done prior to taking the course)d=(number of sit-ups that can be done after taking the course)−(number of sit-ups that can be done prior to taking the course). Use a significance level of α=0.05 for the test. Assume that the numbers of sit-ups are normally distributed for the population both before and after taking the fitness course.

Sit-ups before

24

27

52

31

24

57

28

31

56

51

Sit-ups after

38

42

56

38

42

59

41

37

58

57

Step 1 of 5:State the null and alternative hypotheses for the test.

Step 2 of 5: Find the value of the standard deviation of the paired differences. Round your answer to two decimal places.

Step 3 of 5:Compute the value of the test statistic. Round your answer to three decimal places.

Step 4 of 5: Find the p-value for the hypothesis test. Round your answer to four decimal places.

Step 5 of 5: Draw a conclusion for the hypothesis test.
Answer #1

Given that,
null, H0: Ud = 0
alternate, H1: Ud != 0
level of significance, α = 0.05
from standard normal table, two tailed t α/2 =2.262
since our test is two-tailed
reject Ho, if to < -2.262 OR if to > 2.262
we use Test Statistic
to= d/ (S/√n)
where
value of S^2 = [ ∑ di^2 – ( ∑ di )^2 / n ] / ( n-1 ) )
d = ( Xi-Yi)/n) = -8.7
We have d = -8.7
pooled variance = calculate value of Sd= √S^2 = sqrt [ 1059-(-87^2/10 ] / 9 = 5.794
to = d/ (S/√n) = -4.749
critical Value
the value of |t α| with n-1 = 9 d.f is 2.262
we got |t o| = 4.749 & |t α| =2.262
make Decision
hence Value of | to | > | t α| and here we reject Ho
p-value :two tailed ( double the one tail ) - Ha : ( p != -4.7486 ) = 0.001
hence value of p0.05 > 0.001,here we reject Ho
ANSWERS
---------------
null, H0: Ud = 0
alternate, H1: Ud != 0
test statistic: -4.749
critical value: reject Ho, if to < -2.262 OR if to > 2.262
decision: Reject Ho
p-value: 0.001
we have enough evidence to support the claim that he effect of a physical fitness course on one's physical ability, the number of sit-ups that a person could do in one minute,
both before and after the course, was recorded.